OpenStudy (anonymous):

Which polygon has an interior measure of 900°?

OpenStudy (anonymous):

after a triangle you just add 180 degrees for each side you add

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

use the formulae \[\theta=180(n-2) \] which is similar as mentioned above

OpenStudy (anonymous):

ok cause i needed help

OpenStudy (cwrw238):

so solve the equation for n : 900 = 180(n - 2) and you have it

OpenStudy (anonymous):

thxs salmon

OpenStudy (anonymous):

one more question

OpenStudy (anonymous):

what is 1,298 - 987

OpenStudy (anonymous):

311

OpenStudy (anonymous):

thxs ur so smart